OpenStudy (nuccioreggie):

Which two temperatures have a sum of 0 degrees Celsius? 4 degrees Celsius and 4 degrees Celsius 4 degrees Celsius and −4 degrees Celsius −4 degrees Celsius and −4 degrees Celsius 4 degrees Celsius and 0 degrees Celsius

OpenStudy (nuccioreggie):

@acxbox22

OpenStudy (fibonaccichick666):

again, what have you tried, work?

OpenStudy (nuccioreggie):

?

OpenStudy (littlebird):

With these types of problems, it is difficult to show work. In school, children are often taught to memorize the fact that 4+4=8 (a is wrong) Same with 4+0=4 (d is wrong) Anyway, one could show their work by drawing a number line...but in the future doing that every time you want to subtract some simple numbers is going to be a hassle. Cut him some slack. I have left two options to choose from.

OpenStudy (anonymous):

hmmm

OpenStudy (nuccioreggie):

b

OpenStudy (littlebird):

You are correct

OpenStudy (fibonaccichick666):

work is simple here little bird, you say well, 4+4=? 4+(-4)=? -4-4=? that type of thing, and should be able to explain that reasoning for the work

OpenStudy (fibonaccichick666):

you should not memorize either, but be able to simplify to something if you do not know the answer